OpenStudy (anonymous):

the everton college store paid $1830 for an order of 50 calculators. The store paid $9 for each scientific calculator, the others all graphing calculators cost the store $55 each. How many of each type of calculator was ordered?

OpenStudy (anonymous):

30 graphing, 20 scientific

OpenStudy (anonymous):

The equation is: 1830 = 55x + (50-x)*9 where x = no. of graphing calculators

OpenStudy (anonymous):

9x+55y=1830-(1) , x+y=50-(2), => y=50-x put this in eq (1) then 9x+55(50-x)=1830 => 2750-46x=1830 =>46x=920 =>46x=920 =>x=20 so y=30 ans is 20 scientific calculators & 30graphing calcs